How much force must be applied on a blade of length 4cm and thickness of 0.1mm to exert a pressure of 4000000pa?
1
Expert's answer
2020-08-06T16:46:16-0400

The pressure by definition

"P=\frac{F}{A}"

Hence, the force applied on a blade

"F=PA\\
=(4000000\:\rm Pa)\times(0.04\: m\times 0.0001\: m)=16\: N"